### v1

# TKG-Thinker: Towards Dynamic Reasoning over Temporal Knowledge Graphs via Agentic Reinforcement Learning

Abstract
Temporal knowledge graph question answering (TKGQA) aims to answer time-sensitive questions by leveraging temporal knowledge bases. While Large Language Models (LLMs) demonstrate significant potential in TKGQA, current prompting strategies constrain their efficacy in two primary ways. First, they are prone to reasoning hallucinations under complex temporal constraints. Second, static prompting limits model autonomy and generalization, as it lack optimization through dynamic interaction with temporal knowledge graphs (TKGs) environments. To address these limitations, we propose TKG-Thinker, a novel agent equipped with autonomous planning and adaptive retrieval capabilities for reasoning over TKGs. Specifically, TKG-Thinker performs in-depth temporal reasoning through dynamic multi-turn interactions with TKGs via a dual-training strategy. We first apply Supervised Fine-Tuning (SFT) with chain-of thought data to instill core planning capabilities, followed by a Reinforcement Learning (RL) stage that leverages multi-dimensional rewards to refine reasoning policies under intricate temporal constraints. Experimental results on benchmark datasets with three open-source LLMs show that TKG-Thinker achieves state-of-the-art performance and exhibits strong generalization across complex TKGQA settings.
TKG-Thinker: Towards Dynamic Reasoning over Temporal Knowledge Graphs via Agentic Reinforcement Learning
Zihao Jiang1, Miao Peng3thanks:   Equal contribution., Zhenyan Shan2, Wenjie Xu1, Ben Liu1, Gong Chen1, Ziqi Gao4, Min Peng2 1School of Computer Science, Wuhan University, China 2School of Artificial Intelligence, Wuhan University 3The Hong Kong University of Science and Technology (Guangzhou) 4Tsinghua Shenzhen International Graduate School, Tsinghua University {jiangzihao,bbcavendish,vingerxu,liuben123,chengongcg,pengm}@whu.edu.cn mpeng885@connect.hkust-gz.edu.cn ziqigao@sz.tsinghua.edu.cn
1 Introduction
Temporal knowledge graphs (TKGs) organize factual knowledge over time and serve as an essential foundation for a wide range of knowledge-driven applications, such as recommendation systems Li et al. (2025c); Chen et al. (2025) and question answering Liu et al. (2025b); Gao et al. (2024). In TKGs, facts are represented as quadruples (subject, relation, object, timestamp). Building upon this representation, temporal knowledge graph question answering (TKGQA) focuses on answering time-sensitive questions by leveraging the knowledge stored in TKGs. For instance, the question “Which team did Luka Dončić play for on 2025-02-03?” can be answered using the quadruple (Luka Dončić, play for, Los Angeles Lakers, 2025-02-03).
Figure 1: Comparison between TKG-Thinker and existing LLM-based methods. TKG-Thinker employs a think–action–observation loop for autonomous interaction with TKGs, enabling verified temporal reasoning.
Recently, large language models (LLMs) have demonstrated remarkable performance in tackling complex tasks DeepSeek-AI (2024); Yang et al. (2025); Peng et al. (2025). Building on this success, recent studies have increasingly focused on exploring the potential of LLMs for addressing TKGQA. For instance, some methods QianyiHu et al. (2025); Chen et al. (2024b) employ few-shot prompting to guide LLM-based agents in performing temporal reasoning over TKGs, while others Gong et al. (2025); Qian et al. (2025, 2024) decompose temporal questions into a sequence of sub-questions and combine retrieval-augmented generation (RAG) mechanisms to support step-by-step reasoning. Despite substantial progress, both paradigms struggle in complex settings for two main reasons.
First, existing LLM-based methods are prone to reasoning hallucinations when handling complex temporal constraints in TKGs, including incorrect sub-question decomposition and insensitivity to fine-grained temporal constraints. As illustrated in Figure 1(a), when faced with the question “Before the Ministry of Finance Economy Commerce Trade of South Africa, which country last wanted to negotiate with Yemen?”, LLMs often fail to account for critical temporal constraints such as before and last. As a result, even when the relevant evidence is explicitly available in the temporal context, these methods tend to generate logically inconsistent reasoning steps and ultimately produce incorrect answers. Second, current methods suffer from limited autonomy and suboptimal generalization due to their reliance on static, manually-engineered workflows. More fundamentally, these models lack optimization through dynamic interaction with TKG environments, hindering the development of grounded temporal reasoning. As shown in Figure 1(b), this limitation leads to two critical failures: (1) retrievers often provide context that violates temporal constraints, and (2) the lack of internal verification mechanisms prevents LLMs from autonomously detecting and correcting such misaligned evidence during inference.
Regarding the limitations of static workflows, Reinforcement Learning (RL) Shang et al. (2025); Yue et al. (2025) provides a promising paradigm for shifting to autonomous optimization. By utilizing dynamic reward signals Shao et al. (2024a), RL allows LLMs to acquire complex reasoning behaviors, such as self-correction and strategic search, which are essential for navigating temporal environments Jin et al. (2025). Such emergent capabilities provide a robust mechanism to bridge the gap between static retrieval and temporally-grounded reasoning over TKGs, thereby facilitating the mitigation of hallucinations and limited exploration. Motivated by this perspective, we pose the following research question to guide our study: Can LLMs be effectively trained to autonomously perform dynamic reasoning and retrieval in complex TKGQA scenarios via RL-based optimization?
To address these challenges, we propose TKG-Thinker, a novel agent that reformulates TKGQA as a multi-step interactive process within a dynamic environment. TKG-Thinker performs principled question decomposition and temporal analysis through a two-stage optimization. Specifically, we first employ supervised fine-tuning on a customized dataset with chain-of-thought reasoning paths to equip the model with planning and ReAct-style capabilities Yao et al. (2023). This stage establishes the fundamental "think–action–observation" loop and effectively alleviates the cold-start problem for subsequent optimization. In the second stage, we optimize TKG-Thinker via RL, formalizing temporal reasoning as a sequential decision-making process. To ensure effective policy optimization, we implement a structured interaction protocol where the agent must provide explicit reasoning steps before executing predefined temporal actions (e.g., planning, time-aware retrieval), ensuring that trajectories are fully observable. Specifically, we employ a multi-objective reward mechanism that incorporates an outcome reward for factual correctness, a format reward for structured reasoning, and a retrieval reward for information coverage. This scheme enables TKG-Thinker to internalize autonomous and dynamic reasoning behaviors in complex TKGQA scenarios. In summary, the contributions of this paper are as follows:
  • We introduce TKG-Thinker, a novel agent capable of autonomously performing dynamic, multi-step temporal reasoning.
  • To the best of our knowledge, this is the first work to explore modeling TKGQA as an RL-driven interleaved decision-making process with a time-aware interaction protocol and multi-dimensional reward design.
  • Extensive experiments on benchmark datasets with three open-source LLMs demonstrate significant improvements over state-of-the-art TKGQA methods across multiple metrics.
2 Related Work
2.1 TKGQA
Temporal Knowledge Graph Question Answering is a challenging task that requires models to jointly reason over entities and temporal information in TKGs. Early approaches, such as: MultiQA Chen et al. (2023), TempoQR Mavromatis et al. (2022), and TSQA Shang et al. (2022), typically formulate TKGQA as a temporal knowledge graph completion task, relying on scoring functions to assess the plausibility of candidate facts. Recent LLM-based methods mainly treat the question as a query over the TKGs and use retrieved evidence for reasoning. Specifically, ARI Chen et al. (2024b) enhances the temporal adaptability of LLMs through time-aware training and reasoning signals, while TempAgent QianyiHu et al. (2025) treats the LLM as an agent that performs interaction. TimeR4 Qian et al. (2024) and PoK Qian et al. (2025) strengthen LLM reasoning by improving the retrieval component, whereas RTQA Gong et al. (2025) decomposes questions into sub-problems solved in a bottom-up manner with LLMs and TKGs. Nevertheless, these LLM-based methods still rely on manually crafted prompts, which limits their ability to autonomously detect and correct evidence.
2.2 Agentic Reasoning with RL
While prompting-based methods facilitate search capabilities in a training-free manner Li et al. (2025b), the landscape is shifting toward training-centric agentic reasoning. Advanced approaches have demonstrated that Reinforcement Learning with Verifiable Rewards (RLVR) can unlock superior reasoning abilities in LLMs Shao et al. (2024b); Sheng et al. (2024). This has catalyzed efforts to optimize agentic workflows—such as multi-step search Jin et al. (2025); Team et al. (2025) and external tool integration Shang et al. (2025)—using RL. Despite progress in long-horizon tasks through self-reflection Shi et al. (2025) and structured memory Yan et al. (2025), these methods primarily focus on logical or mathematical tasks, but lack specialized mechanisms to navigate the intricate temporal constraints inherent in TKGs.
3 Preliminary
To enhance the reasoning capabilities of LLMs for TKGQA, we employ the RLVR framework, which optimizes the policy $\pi_{\theta}$ using deterministic rewards $r$ (e.g., execution results or exact-match accuracy). Formally, the objective is to maximize:
$\mathcal{J}(\theta)$ $=\hat{\mathbb{E}}_{Q,\mathbf{y}\sim\pi_{old}}\left[\frac{1}{G}\sum_{i=1}^{G}f_{\epsilon}(\rho_{i}(\theta),\hat{A}_{i})\right]$
$-\beta\cdot\hat{\mathbb{E}}_{Q}\left[\mathbb{D}_{KL}[\pi_{\theta}(\cdot|Q)||\pi_{\text{ref}}(\cdot|Q)]\right],$
where $G$ denotes the number of sampled trajectories per prompt ($G>1$ for GRPO). $\rho_{i}(\theta)=\frac{\pi_{\theta}(y_{i}|Q)}{\pi_{old}(y_{i}|Q)}$ represents the importance sampling ratio. $f_{\epsilon}$ denotes the clipping function used in PPO/GRPO to stabilize updates, while $\hat{A}_{i}$ represents the advantage of trajectory $y_{i}$ computed based on verifiable rewards. The KL divergence term, scaled by $\beta$, penalizes deviations from a reference policy $\pi_{\text{ref}}$ to prevent model collapse.
4 Methodology
Figure 2: The overview of our proposed TKG-Thinker. We first apply supervised fine-tuning on high-quality trajectories to mitigate the cold-start problem, and further refine the model via online reinforcement learning with temporal tool calls. The bottom panel illustrates three rollouts: a complete success, a partial success, and a failure.
In this section, we introduce TKG-Thinker, a novel agent equipped with autonomous planning and adaptive retrieval capabilities for reasoning over TKGs. As illustrated in Figure 2, TKG-Thinker is trained through two complementary stages: (1) Supervised Fine-Tuning (SFT) for cold start (§ 4.1), and (2) Online Reinforcement Learning with Temporal Tool Calls (§ 4.2).
4.1 Supervised Fine-Tuning for Cold Start
TKG-Thinker relies on meaningful exploration over tool-augmented trajectories, but a generic base model does not yet know how to plan or invoke tools in the expected format, leading to low-quality rollouts Shao et al. (2025). To address this, we perform SFT on trajectories generated by a strong teacher model (e.g., GPT-4o) acting as a tool-augmented agent Li et al. (2025a), thereby initializing TKG-Thinker with a reasonable search and citation strategy before online RL. To enable the teacher model to produce tool-augmented CoT trajectories suitable for SFT, we carefully construct a prompting pipeline. Specifically, we first adopt a few-shot prompting strategy to elicit structured CoT trajectories from the teacher model, as detailed in Appendix A.1. However, not all generated trajectories are reliable. Therefore, we apply a two-stage rejection sampling pipeline:
  • Format validity filtering: We discard trajectories that violate structural constraints, ensuring consistent CoT patterns.
  • Answer correctness filtering: We filter out trajectories whose final answer does not match the ground-truth label in the training set.
In this way, we final curated set forms a high-quality CoT dataset for reasoning activation, and statistical details are provided in Appendix A.3. With this dataset in hand, given a question $Q$ and a filtered trajectory $y=[y_{1},\dots,y_{T}]$, where $y$ concatenates the structured reasoning steps, tool calls, and final answer, the SFT objective maximizes the likelihood of the teacher trajectory:
$\mathcal{L}_{\mathrm{SFT}}=-\sum_{t=1}^{T}\log\pi_{\theta}(y_{t}\mid Q,y_{<t}),$ (1)
where $\pi_{\theta}$ denotes the model’s token distribution. Following prior work Li et al. (2025a), we compute the loss only over model-generated tokens and exclude environment feedback (e.g., observations). The resulting model $\pi_{\mathrm{SFT}}$ serves as the initialization for the second online RL stage, substantially alleviating cold-start issues.
4.2 Online RL with Temporal Tools Call
After supervised fine-tuning stage, we further apply the online reinforcement learning with temporal tool calls, which enhances its multi-hop and time-sensitive reasoning capabilities. To enable effective learning in this setting, we address the problem from three key perspectives: action space, reward design, and training objective.
4.2.1 Action Space
As we can observe the phenomenon that existing retrievers are insensitive to temporal constraints and LLMs tend to overlook temporal requirements and hallucinate during question decomposition Qian et al. (2025); Guo et al. (2025), we identify the critical reasoning characteristics of temporal questions, and we categorize the reasoning paradigms of temporal reasoning over TKGs into specific taxonomies according to timestamps. Crucially, we formalize the agent’s action space as a collection of temporal functional tools. By doing so, we transform temporal reasoning into actionable primitives that directly mirror the TKG structure, moving beyond the limitations of the LLM’s implicit reasoning capabilities. In detail, our action space includes think, plan, temporal search actions, and answer, with the temporal search actions defined as follows:
  • Search_time(query). This action returns timestamps or time intervals associated with relevant quadruples.
  • Search_specific(query, $t$). This action returns relevant quadruples at the specified time $t$.
  • Search_before(query, $t$). This action returns relevant quadruples occurring strictly before the specified time $t$.
  • Search_after(query, $t$). This action returns relevant quadruples occurring strictly after the specified time $t$.
  • Search_between(query, $t_{1},t_{2}$). This action returns relevant quadruples occurring within the time interval $t_{1},t_{2}]$.
Based on this action space and treating TKGs as the environment, temporal search tool returns structured feedback enclosed within <observation> and </observation>. Under this formulation, we adopt a ReAct-style Yao et al. (2023); Liu et al. (2025a) interaction protocol, in which the reasoning process proceeds as a a sequence of planning, internal thought generation, temporal search tool calls, and environment feedback. Formally, the interaction trajectory at step $n$ can be represented as:
$\mathcal{H}_{n}=(\tau_{0},p,\tau_{1},a_{1},o_{1},\ldots,\tau_{n},a_{n-1},o_{n-1}),$ (2)
where $\tau$ denotes the agent’s internal thought, $a$ is an action selected from the temporal tool set, with $p$ being the initial planning action, and $o$ is the observation obtained by executing the action $a$ over the TKGs. Based on the historical trajectory $\mathcal{H}_{n}$, the generation process for the next thought $\tau_{n}$ and action $a_{n}$ can be formulated as:
$\pi_{\theta}(\tau_{n}\mid\mathcal{H}_{n})=\prod_{i=1}^{|\tau_{n}|}\pi_{\theta}(\tau_{n}^{i}\mid\mathcal{H}_{n},\tau_{n}^{<i}),$ (3)
$\pi_{\theta}(a_{n}\mid\mathcal{H}_{n},\tau_{n})=\prod_{j=1}^{|a_{n}|}\pi_{\theta}(a_{n}^{j}\mid\mathcal{H}_{n},\tau_{n},a_{n}^{<j}),$ (4)
where $\pi_{\theta}=\pi_{\mathrm{SFT}}$, $\tau_{n}^{i}$ and $|\tau_{n}|$ denote the $i$-th token and the length of $\tau_{n}$, and $a_{n}^{j}$ and $|a_{n}|$ denote the $j$-th token and the length of $a_{n}$. The interaction loop terminates when either the answer action is invoked or the interaction-turn budget $B_{\max}$ is reached.
4.2.2 Reward Design
To optimize the above interactive process, we adopt the RLVR framework equipped with a novel multi-reward formulation. We incorporate three key components into our reward design: the format reward, the retrieval reward, and the outcome reward.
Format Reward verifies whether the generated rollout adheres to the structured interaction protocol. Specifically, we ensure that the entire rollout follows the temporal interaction pattern defined in Eq. 2. Formally, the format reward is defined as:
$R_{\mathrm{fmt}}=\alpha\,\mathbb{I}_{\mathrm{fmt}},$ (5)
where $\mathbb{I}_{\mathrm{fmt}}\in\{0,1\}$ is a binary format validity indicator and $\alpha\in(0,1)$ is a scaling coefficient.
Retrieval Reward measures whether the retriever successfully retrieves evidence containing the correct answer, defined as:
$R_{\mathrm{ret}}=\gamma\,\mathbb{I}_{\mathrm{ret}},$ (6)
where $\mathbb{I}_{\mathrm{ret}}\in\{0,1\}$ is a binary retrieval indicator, and $\gamma\in(0,1)$ is a tunable scaling coefficient.
Outcome Reward evaluates the correctness of final output answer $a_{\mathrm{pred}}$ by comparing it against the ground truth answer $a_{\mathrm{gold}}$ using rule-based criteria exact match (EM):
$R_{\mathrm{out}}=\mathrm{EM}(a_{\mathrm{pred}},a_{\mathrm{gold}}),$ (7)
where $\mathrm{EM}(\cdot,\cdot)$ returns $1$ if the two strings match exactly and 0 otherwise, and thus $R_{\mathrm{out}}\in\{0,1\}$. Finally, we combine the above components into the overall reward. The final reward $R_{\mathrm{all}}$ is defined as:
$R_{\mathrm{all}}={}$ $R_{\mathrm{out}}\!\left(1-(1-\mathbb{I}_{\mathrm{fmt}})\lambda\right)$ (8)
$+(1-R_{\mathrm{out}})\!\left(R_{\mathrm{fmt}}+R_{\mathrm{ret}}\right)$
$+(1-R_{\mathrm{out}})\,\delta\left(1-\mathbb{I}_{\mathrm{fmt}}\right),$
where $\lambda>0$ denotes the penalty applied when the answer is correct but the format is invalid, and $\delta>0$ serves as a fallback reward granted when both the answer and format are incorrect. In this way, $R_{\mathrm{fmt}}$ enforces adherence to the temporal interaction protocol, $R_{\mathrm{ret}}$ promotes effective problem decomposition and evidence retrieval over TKGs, and $R_{\mathrm{out}}$ ensures factual correctness.
4.2.3 Training Objective
With the multi-dimensional reward formulation defined, we formalize the overall training objective of TKG-Thinker framework. To ensure robust policy optimization, we adopt the RLVR paradigm, which can be instantiated through either PPO or GRPO. The policy $\pi_{\theta}$ is optimized by maximizing the objective function $\mathcal{J}(\theta)$, which encourages trajectories with higher-than-average rewards while maintaining stability via importance sampling and Kullback–Leibler (KL) divergence constraints. The overall objective is defined as:
[…]
$-\beta\cdot\hat{\mathbb{E}}_{Q}\left[\mathbb{D}_{KL}\left[\pi_{\theta}(\cdot|Q)||\pi_{\text{ref}}(\cdot|Q)\right]\right],$
[…]
5 Experiments
In this section, we evaluate TKG-Thinker on widely used datasets. We conduct extensive experiments to demonstrate the effectiveness of our method by answering the following research questions (RQ): (1) RQ1: How does TKG-Thinker perform compared to state-of-the-art baselines on complex TKGQA datasets? (2) RQ2: What is the contribution of each key module in the TKG-Thinker framework to the overall performance? (3) RQ3: How do different retrieval configurations affect reasoning performance? (4) RQ4: How does RL optimization shape the model’s behavior in TKGQA scenarios? We also conduct a cross-domain generalization study in Appendix B, and present a case study in Appendix E to further demonstrate the robustness and the advantages of our proposed method.
5.1 Experimental Settings
We evaluate TKG-Thinker on representative TKGQA benchmarks, including MULTITQ Chen et al. (2023) and CronQuestions Saxena et al. (2021). Detailed descriptions of these datasets are provided in Appendix A.2, while additional results on the cross-domain benchmark TimelineKGQA Sun et al. (2025) are reported in Appendix B. We adopt Hits@1 as the evaluation metric, measuring the proportion of questions for which the top-ranked prediction is correct. We compare TKG-Thinker against three categories of baselines: PLM-based methods, Embedding-based methods, and LLM-based methods. A detailed description of all baseline models is provided in Appendix A.4. For training, we use GPT-4o as the teacher model to generate trajectories and adopt e5-base-v2 Wang et al. (2022) for evidence retrieval, retrieving the top-15 most relevant quadruples per query. TKG-Thinker is instantiated with Llama3-8B-Instruct, Qwen2.5-7B-Instruct, and Qwen3-4B-Instruct-2507 as backbone models. Additional implementation details are provided in Appendix A.5.
MULTITQ CronQuestions
Methods Overall Question Type Answer Type Overall Question Type Answer Type
Single Multiple Entity Time Simple Complex Entity Time
[…]
BERT Devlin et al. (2019) 0.083 0.092 0.061 0.101 0.040 0.243 0.249 0.239 0.277 0.179
DistillBERT Sanh et al. (2019) 0.083 0.087 0.074 0.102 0.037
ALBERT Lan et al. (2020) 0.108 0.116 0.086 0.139 0.032 0.248 0.255 0.235 0.279 0.177
[…]
EmbedKGQA Saxena et al. (2020) 0.206 0.235 0.134 0.290 0.001 0.288 0.290 0.286 0.411 0.057
CronKGQA Saxena et al. (2021) 0.279 0.337 0.134 0.328 0.156 0.647 0.987 0.392 0.699 0.549
MultiQA Chen et al. (2023) 0.293 0.347 0.159 0.349 0.157
[…]
ARI Chen et al. (2024b) 0.380 0.680 0.210 0.394 0.344 0.707 0.860 0.570 0.660 0.800
Naive RAG Chen et al. (2024a) 0.379 0.469 0.155 0.242 0.672 0.633 0.726 0.280 0.610 0.684
ReAct RAG Yao et al. (2023) 0.398 0.506 0.130 0.243 0.735 0.809 0.863 0.600 0.768 0.895
TempAgent QianyiHu et al. (2025) 0.702 0.857 0.316 0.624 0.870 0.842 0.895 0.640 0.805 0.921
TimeR4 Qian et al. (2024) 0.728 0.887 0.335 0.639 0.945
RTQA Gong et al. (2025) 0.765 0.902 0.424 0.692 0.942
PoK Qian et al. (2025) 0.779 0.929 0.409 0.696 0.962
[…]
♣  Qwen2.5-7B-Instruct 0.855 0.910 0.721 0.814 0.955 0.893 0.958 0.844 0.863 0.949
♠  Qwen2.5-7B-Instruct 0.824 0.881 0.683 0.774 0.945 0.868 0.937 0.817 0.838 0.925
♣  Llama3-8B-Instruct 0.810 0.872 0.659 0.756 0.943 0.915 0.968 0.875 0.893 0.955
♠  Llama3-8B-Instruct 0.799 0.861 0.644 0.709 0.739 0.873 0.949 0.817 0.844 0.928
♣ Qwen3-4B-Instruct 0.804 0.914 0.533 0.739 0.963 0.889 0.953 0.837 0.857 0.948
♠  Qwen3-4B-Instruct 0.780 0.901 0.480 0.708 0.956 0.877 0.953 0.820 0.841 0.944
Table 1: Performance comparison of baselines and TKG-Thinker in Hits@1 across different question and answer types on MULTITQ and CronQuestions. ♣ denotes TKG-Thinker trained with SFT+GRPO, while ♠ denotes training with SFT+PPO. The best and second-best scores are marked in bold and underline, respectively.
5.2 Main Results (RQ1)
In this section, we compare TKG-Thinker with representative baselines on MULTITQ and CronQuestions. As shown in Table 5.1, TKG-Thinker achieves consistently superior overall performance, outperforming diverse baselines across different model families, parameter scales (4B, 7B, 8B), and RL training strategies (trained with GRPO or PPO), demonstrating strong model-agnostic applicability. Compared to the strongest baseline, TKG-Thinker achieves absolute overall Hits@1 improvements of $7.60\%$ and $7.30\%$ on MULTITQ and CronQuestions, respectively. These results suggest that enabling LLMs to dynamically interact with TKGs via RAG mechanisms facilitates effective search strategies and temporally grounded reasoning capabilities. Notably, TKG-Thinker exhibits substantial improvements on complex multi-step TKGQA tasks, surpassing the best-performing baselines on the corresponding complex settings by $29.70\%$ on MULTITQ (Multiple) and $23.50\%$ on CronQuestions (Complex). This further confirms that our approach significantly enhances temporal multi-hop reasoning through explicit planning and time-aware retrieval tool usage. While TKG-Thinker shows slightly lower performance on Single-type questions in MULTITQ, this difference can be reasonably attributed to PoK’s use of a retriever specifically optimized for single-step temporal retrieval.
Model Overall Answer Type Fine-grained Question Type
Entity Time Equal Before/After First/Last Equal Multi After First Before Last
TKG-Thinker 0.855 0.814 0.955 0.939 0.851 0.920 0.792 0.678 0.729
w/o SFT Stage 0.591 0.465 0.897 0.869 0.609 0.836 0.213 0.078 0.084
w/o Plan Action 0.796 0.766 0.871 0.933 0.822 0.838 0.704 0.538 0.609
w/o Temporal Retrievers 0.458 0.252 0.959 0.466 0.456 0.909 0.205 0.133 0.139
Table 2: Ablation study on the MULTITQ dataset. Bold indicates the best performance, while underline marks the second-best . Single-type questions include Equal, Before/After, and First/Last; Multiple-type questions include Equal Multi, After First, and Before Last. “w/o” means removing or replacing the corresponding module.
Figure 3: Retriever analysis on the MULTITQ dataset. Left: Performance comparison of different retriever models. Right: Effect of retrieval depth, measured by the number of top-$k$ retrieved quadruples.
Figure 4: Training dynamics of TKG-Thinker implemented with GRPO and PPO on MULTITQ. Left: Training Reward; Middle: Retrieval Call Steps; Right: Action Steps.
5.3 Ablation Study (RQ2)
In this section, we conduct a series of ablation experiments to examine the contribution of each component in TKG-Thinker, as summarized in Table 2, including the SFT stage, the planning mechanism, and the temporal retrievers. Specifically, we systematically remove or replace individual components to construct corresponding model variants for comparison. Additional ablation results on the CronQuestions are reported in Appendix C.
Effect of the SFT Stage. We further analyze the role of SFT, which initializes the model’s ability to execute structured reasoning protocols prior to reinforcement learning. When the SFT stage is removed and the model is trained with RL alone, the overall performance drops drastically by $26.40\%$. This confirms that SFT provides essential scaffolding that stabilizes learning, reduces temporal hallucination, enables verifiable temporal reasoning rather than unconstrained free-form generation.
Effect of the Plan Action. We remove the Plan action and prompt the model to interact directly with the environment using the original queries. As a result, eliminating the planning component leads to an overall performance drop of 5.90%. In particular, performance on Multiple-type temporal questions decreases by 8.80%, 14.00%, and 12.00% on Equal Multi, After First, and Before Last, respectively. These results suggest that planning component plays an important role in reliable temporal reasoning over TKGs, as removing it tends to hallucinated intermediate reasoning steps.
Effect of the Temporal Retrievers. To assess the role of temporal retrievers, we replace them with a purely semantic retriever that ignores temporal constraints. This substitution yields the largest performance degradation ($39.70\%$), highlighting the importance of temporal alignment between questions and evidence. Notably, performance drops sharply on multiple-type temporal questions, demonstrating that temporal retrieval is indispensable for providing fine-grained, temporally grounded evidence to support reliable temporal reasoning.
5.4 Retrieval Analysis (RQ3)
Effect of Retriever Model. Since retrieval quality directly determines the availability of temporal evidence, we first evaluate TKG-Thinker under four representative retrievers: e5-base-v2, bge-m3, contriever, and qwen3-embedding-4B. As shown in Figure 3 (Left), all four retrievers achieve competitive performance, substantially surpassing the strongest baseline (PoK). Among them, contrastively trained retrievers (e.g., e5-base-v2, bge-m3) deliver superior performance, with the advantage being more evident on Multiple questions.
Effect of Retrieval Depth. The hyperparameter $k$ controls how many top-ranked quadruples the temporal search tools return as environmental feedback. As illustrated in Figure 3 (Right), performance increases with $k$ and then declines. This reflects a trade-off: larger $k$ improves the likelihood of retrieving useful evidence, whereas excessively large $k$ introduces distractors that impede LLM reasoning. Notably, performance degradation at larger $k$ is more pronounced on Multiple questions. We attribute this phenomenon to the accumulation of errors across successive reasoning steps, as Multiple-type questions require iterative retrieval and multi-step reasoning, where early errors are progressively amplified in later stages. In practice, we find that $k=15$ offers the best balance between evidence coverage and distractor noise.
5.5 Training Dynamics (RQ4)
To investigate how TKG-Thinker evolves during the training process, we illustrate its training dynamics in Figure 4 (with additional entropy and response details in Appendix D). As shown in the left panel, both PPO and GRPO exhibit a steady increase in training rewards. This demonstrates that our fine-grained reward design provides stable reinforcement signals, facilitating consistent policy optimization. Regarding action and retrieval dynamics, we observe a clear "decrease–then–increase" pattern. Specifically, the average number of action steps initially drops sharply as the model learns to follow the required output format and eliminates redundant or invalid actions. Subsequently, both action steps and retrieval calls gradually increase and stabilize, indicating that TKG-Thinker strategically invokes additional temporal tool calls to acquire necessary evidence and thereby strengthens its agentic reasoning capability. Notably, while both algorithms converge well, PPO achieves a higher reward ceiling and more frequent retrieval calls in the later stages of training.
6 Conclusion
we introduce TKG-Thinker, a novel agent equipped with autonomous planning and adaptive retrieval capabilities for reasoning over TKGs. By modeling the TKG as a dynamic environment, TKG-Thinker integrates supervised fine-tuning and reinforcement learning with a multi-reward optimization scheme to enhance temporal reasoning. Experiments show that TKG-Thinker consistently outperforms baselines, demonstrating the effectiveness of explicit interaction and RL-driven optimization in reducing hallucination and improving multi-step reasoning.
Limitations
Despite TKG-Thinker’s strong performance achieved in TKGQA, this work remains several limitations. The current reward mechanism relies heavily on binary indicators and rule-based criteria, such as Exact Match (EM) for outcomes and basic format verification. This outcome-based reward lacks a nuanced evaluation of the intermediate reasoning process. Future iterations could incorporate an LLM Judge with detailed rubrics to qualitatively assess the logical and temporal consistency of the think and plan steps, ensuring that the model understands complex temporal constraints rather than just optimizing for a specific output format. Besides, while the model demonstrates effective multi-step reasoning, the relative simplicity of current datasets and benchmarks—particularly their limited reasoning hops—restricts the training of temporal agents capable of long-range planning and inference. Future work should thus explore more complex synthetic multi-hop tasks and open-world settings to foster greater model robustness.
Ethics Statement
In constructing the CoT-based SFT datasets, we have taken into account ethical considerations and limitations commonly associated with large language models. All data used in this work are publicly available and do not contain personal or sensitive information. Nonetheless, we acknowledge that, despite our best efforts, the datasets may still contain gaps or unintended biases. To mitigate these concerns, the source data has been curated to ensure diversity and reduce potential bias. Through careful dataset construction, review, and testing procedures, we strive to uphold ethical AI principles while advancing research in TKGQA.
Appendix A Experimental Settings
A.1 Few-shot Prompt for Initial Trajectory Generation
As shown in Figure 5, we adopt a few-shot prompting strategy to elicit structured, tool-augmented reasoning during trajectory generation. This stage provides the initial pool of trajectories before subsequent format and answer-correctness filtering.
Figure 5: Few-shot Prompt for Generating Trajectories.
A.2 Dataset Details
Type Train Valid Test
Single Equal 135,890 18,983 17,311
Before/After 75,340 11,655 11,073
First/Last 72,252 11,097 10,480
Multiple Equal Multi 16,893 3,213 3,207
After First 43,305 6,499 6,266
Before Last 43,107 6,532 6,247
Total 386,787 57,979 54,584
Table 3: Data statistics of MULTITQ.
Type Train Valid Test
Simple Entity 90,651 7,745 7,812
Simple Time 61,471 5,197 5,046
Before/After 23,869 1,982 2,151
First/Last 118,556 11,198 11,159
Time Join 55,453 3,878 3,832
Simple Reasoning 152,122 12,942 12,858
Complex Reasoning 197,878 17,058 17,142
Entity Answer 225,672 19,362 19,524
Time Answer 124,328 10,638 10,476
Total 350,000 30,000 30,000
Table 4: Dataset Statistics of CronQuestions.
MULTITQ.
MULTITQ is a large-scale temporal question answering dataset that incorporates multi-granularity temporal information. It provides a comprehensive evaluation protocol across several dimensions: Question Type (Multiple vs. Single), Answer Type (Entity vs. Time), and Time Granularity (year, month, and day). The detailed statistics of MULTITQ are presented in Table 3.
CronQuestions.
CronQuestions is a temporal QA benchmark consisting of 410K unique question–answer pairs. Its questions can be categorized into two major types: Simple temporal reasoning (e.g., Simple Entity and Simple Time) and Complex temporal reasoning (e.g., Before/After, First/Last, and Time-Join queries), depending on the temporal constraints involved. Detailed statistics of CronQuestions are shown in Table 4.
A.3 Statistics of the CoT Dataset for SFT
Table 5: Statistics of the SFT datasets constructed from MULTITQ and CronQuestions.
Dataset Question Type Count
MULTITQ Equal 4017
Before/After 1668
First/Last 1062
Equal Multi 1395
After First 1074
Before Last 1055
Total 10271
CronQuestions Simple Entity 2509
Simple Time 1726
Before/After 668
First/Last 3301
Time Join 1508
Total 9712
To mitigate the cold-start issue in reinforcement learning, we construct CoT-style supervised fine-tuning datasets from MULTITQ and CronQuestions. As summarized in Table 5, these datasets cover diverse temporal reasoning types and provide explicit supervision signals for temporal decomposition and retrieval behaviors.
A.4 Baseline Details
we compare TKG-Thinker against three categories of baselines: (1) PLM-based methods, including BERTDevlin et al. (2019), ALBERTLan et al. (2020), and DistilBERTSanh et al. (2019); (2) Embedding-based methods, such as EmbedKGQASaxena et al. (2020), CronKGQASaxena et al. (2021), and MultiQAChen et al. (2023); and (3) LLM-based methods, including Naive RAG Chen et al. (2024a), ReAct RAGYao et al. (2023), ARIChen et al. (2024b), RTQAGong et al. (2025), and TempAgentQianyiHu et al. (2025), TimR4 Qian et al. (2024), and PoK Qian et al. (2025). For consistency with prior work, we adopt the baseline results reported in Qian et al. (2025), Chen et al. (2024b), and QianyiHu et al. (2025) for comparison.
A.5 Implementation Details
Table 6: Key hyperparameters used in model training.
Category Value
Supervised Fine-tuning (SFT)
Finetuning Type Full
Epochs 4
Batch Size 4
Grad. Accumulation 4
Learning Rate $1\times 10^{-5}$
Scheduler Cosine Decay
Warmup Ratio 0.03
Cutoff Length 8192
Precision BF16
Reinforcement Learning (RL)
Algorithm PPO / GRPO
Training Steps 80
Train Batch Size 256
Max Turns ($B_{\max}$) 8
Temperature 0.7
Retriever Top-$k$ 15
Max Response Length (per turn) $512$
Max Observation Length (per turn) $1024$
Format Reward Coef. ($\alpha$) $0.2$
Outcome Penalty Coef. ($\lambda$) $0.4$
Retrieval Reward Coef. ($\gamma$) $0.1$
Fallback Reward Coef. ($\delta$) $0.1$
Inference
Temperature $0.01$
Top-pp $0.95$
Max New Tokens $512$
Max Infer Step 88
As summarized in Table 6, during the SFT stage, we fine-tune the models via the LLaMA-Factory framework Zheng et al. (2024) with a batch size of 4 for 4 epochs using AdamW with a learning rate of $1\mathrm{e}{-5}$ and cosine decay scheduling. In the RL stage, we switch to the Verl framework and train both PPO and GRPO policies with a batch size of 256, a mini-batch size of 32, and 5 rollouts. We set the interaction-turn budget to $B_{\max}=8$. For rollout collection, we apply sampling with a temperature of 0.7 and conduct retrieval-augmented interactions using the top-$15$ evidence candidates per query. To stabilize trajectory generation and prevent excessive growth of reasoning tokens, model-generated responses and retrieved observations are truncated to $512$ and $1024$ tokens per turn, respectively. Regarding the training data, the SFT stage uses rejection sampling to retain trajectories, while the RL stage is trained on TKGQA data not used in SFT (5,001 QA pairs for MULTITQ and 5,230 for CronQuestions). During inference, we disable stochastic sampling and adopt deterministic decoding with temperature $0.01$, and top-$p=0.95$. All experiments are implemented in PyTorch and conducted on 8 NVIDIA A800 (80GB) GPUs.
Appendix B Cross-domain Generalization Study
Dataset Type Train Val Test
Timeline-CronQuestion Simple 7,200 2,400 2,400
Medium 8,252 2,751 2,751
Complex 9,580 3,193 3,193
Total 25,032 8,344 8,344
Timeline-ICEWS Simple 17,982 5,994 5,994
Medium 15,990 5,330 5,330
Complex 19,652 6,550 6,550
Total 53,624 17,874 17,874
Table 7: Data statistics of Timeline-CronQuestion and Timeline-ICEWS.
Model Overall Simple Medium Complex
RAG baseline 0.235 0.704 0.092 0.009
LLaMA2-7B 0.169 0.049 0.143 0.282
GPT-4o 0.206 0.069 0.130 0.376
RTQA 0.298 0.608 0.218 0.135
[…]
Qwen2.5-7B-Instruct 0.460 0.567 0.294 0.523
Llama3-8B-Instruct 0.491 0.612 0.320 0.546
Table 8: Results on the Timeline-CronQuestion dataset across different reasoning difficulty levels. Bold indicates the best performance and underline denotes the second best performance.
Model Overall Simple Medium Complex
RAG baseline 0.265 0.660 0.128 0.011
LLaMA2-7B 0.111 0.035 0.066 0.322
GPT-4o 0.113 0.051 0.035 0.353
[…]
Qwen2.5-7B-Instruct 0.508 0.556 0.409 0.583
Llama3-8B-Instruct 0.462 0.596 0.273 0.533
Table 9: Results on the Timeline-ICEWS dataset across different reasoning difficulty levels. Bold indicates the best performance, while underline denotes the second-best performance.
Model Overall Answer Type Fine-grained Question Type
Entity Time Simple Entity Simple Time Before/After First/Last Time Join
TKG-Thinker 0.893 0.863 0.949 0.946 0.977 0.699 0.891 0.787
w/o SFT Stage 0.638 0.732 0.568 0.861 0.532 0.487 0.564 0.623
w/o Plan Action 0.860 0.811 0.951 0.942 0.977 0.538 0.899 0.603
w/o Temporal Retrievers 0.821 0.756 0.942 0.924 0.963 0.477 0.891 0.414
Table 10: Ablation study on the CronQuestions dataset. Bold indicates the best performance, while underline marks the second-best. Simple-type questions include Simple Entity, and Simple Time; Complex-type questions include Before/After, First/Last, and Time Join. “w/o” means removing or replacing the corresponding module.
We further evaluate the cross-domain generalizability of TKG-Thinker. Specifically, we evaluate TKG-Thinker on the TimelineKGQA Sun et al. (2025) benchmark, which includes Timeline-CronQuestions and Timeline-ICEWS. These datasets differ in both temporal representations and the complexity of temporal reasoning. In TimelineKGQA, Simple questions require a single contextual fact and typically involve temporally constrained retrieval or timeline position identification; Medium questions require two contextual facts and further involve the combination of retrieval with temporal semantic operations and timeline arithmetic; Complex questions require three contextual facts and cover the full spectrum of temporal reasoning capabilities. The dataset statistics are summarized in Table 7. For this evaluation, we use Qwen2.5-7B-Instruct and Llama3-8B-Instruct as backbones, both trained on MULTITQ using SFT and GRPO. Following Gong et al. (2025) and Sun et al. (2025), we compare TKG-Thinker with a RAG baseline, LLaMA2-7B, GPT-4o, and RTQA. For consistency with prior work, baseline results are sourced from Gong et al. (2025).
As shown in Table B and Table B, TKG-Thinker consistently outperforms all baselines in Overall score, and the results reveal three consistent trends. First, TKG-Thinker exhibits clear performance gains as temporal reasoning complexity increases. While improvements on Simple questions are marginal, the gains on Medium and Complex questions are substantially larger, indicating that TKG-Thinker is particularly effective at handling compositional temporal reasoning and timeline arithmetic. Second, the relative advantage of TKG-Thinker is preserved across datasets with distinct temporal characteristics: on Timeline-CronQuestion the improvements are most salient on Medium and Complex queries, whereas on Timeline-ICEWS the model maintains strong performance across all difficulty levels despite its larger event space and more diverse temporal expressions. Third, LLaMA2-7B and GPT-4o exhibit limited ability in settings that require explicit temporal grounding, whereas retrieval-augmented methods that rely on sub-question decomposition, such as RTQA, partially close the performance gap but still struggle with multi-hop temporal composition. These findings suggest that structured tool-augmented trajectories and agentic decision-making are key to improving temporal reasoning capabilities in various TKGQA scenarios.
Figure 6: Training dynamics of TKG-Thinker implemented with GRPO and PPO on MULTITQ. Left: Generation Entropy; Right: Response Length.
Appendix C Ablation Study on CronQuestion
We further conduct ablation experiments on CronQuestions to analyze the contribution of each component in TKG-Thinker, as summarized in Table 10. Consistent with the findings on MULTITQ, several clear observations can be drawn. First, removing the SFT stage and training the model with reinforcement learning alone leads to a substantial performance drop of 25.50% in overall accuracy. This result indicates that SFT is essential for alleviating the cold-start problem in reinforcement learning by providing a stable initialization and structured reasoning priors. Second, eliminating the Plan action leads to a noticeable performance decline of 3.30% in the overall score, with the most severe degradation observed on Complex-type temporal questions. This suggests that explicit planning is particularly important for handling complex temporal dependencies that require multi-step reasoning. Finally, replacing the temporal retrievers with a purely semantic retriever that disregards temporal constraints leads to the performance drop of $7.2\%$. This result highlights that explicit temporal alignment between the question and the retrieved evidence is a fundamental prerequisite for accurate TKGQA. Overall, these results demonstrate that TKG-Thinker benefits from the combined effects of supervised initialization, explicit planning, and temporally aware retrieval to achieve robust temporal reasoning.
Appendix D Training Dynamics Details
Figure 6 illustrates the evolution of training entropy and response length, providing deeper insights into the training process of TKG-Thinker.
Training Entropy.
The left panel shows that after an initial exploration phase, the policy entropy for both PPO and GRPO consistently declines. This trend indicates that the model is successfully converging from diverse exploration to a stable, optimized policy by exploiting high-reward reasoning paths.
Response Length.
The right panel exhibits a distinct "V-shaped" trajectory:
  • Alignment Phase (Steps 0-15): A sharp decline in length occurs as the model learns to prune redundant tokens and strictly follow the required agentic interaction format.
  • Reasoning Expansion (Steps 15-60): The length gradually increases and stabilizes, suggesting that once the format is mastered, TKG-Thinker learns to generate more substantial reasoning chains and necessary tool calls to solve complex temporal queations.
Figure 7: A Case Study on MULTITQ for Equal Questions.
Figure 8: A Case Study on MULTITQ for Equal Multi Questions.
Figure 9: A Case Study on MULTITQ for After First Questions.
Figure 10: A Case Study on MULTITQ for Before Last Questions.
Appendix E Case Study
To illustrate how TKG-Thinker, equipped with autonomous planning and adaptive retrieval, performs temporal reasoning over TKGs to obtain correct answers, we analyze several representative questions.
MULTITQ.
Figures 7 and 8 present case studies of Equal and Equal Multi questions, respectively. These examples demonstrate that TKG-Thinker can accurately identify relevant temporal conditions, retrieves supporting quadruples for each constraint, and integrates the evidence to derive the final answer. Figures 9 and 10 further illustrate After First and Before Last questions. In these cases, TKG-Thinker first plans an anchor event, then performs iterative retrieval under strict temporal constraints, and crucially conducts internal temporal verification to ensure that no earlier or later events violate the query requirements. For example, Figure 10 shows that TKG-Thinker performs bounded verification via repeated Search_between(query, $t_{1},t_{2}$) calls, updating the candidate from Association_of_Southeast_Asian_Nations to Qatar and finally to Japan. These examples provide concrete evidence of the model’s verification behavior during temporal reasoning.
Figure 11: A Case Study on CronQuestions for Simple TKGQA Questions.
Figure 12: A Case Study on CronQuestions for Complex TKGQA Questions.
CronQuestions.
Unlike MULTITQ, which represents facts as timestamped quadruples, CronQuestions models knowledge using quintuples with explicit temporal intervals. As illustrated in Figure 11 and Figure 12, both simple and complex queries in CronQuestions require reasoning over interval-level temporal constraints. In this setting, TKG-Thinker is able to decompose a given question into a sequence of executable subtasks via the plan action, perform iterative and temporally aligned retrieval, and ultimately produce the correct answer.
Figure 13: A Case Study on TimelineKGQA.
TimelineKGQA.
Figure 13 illustrates a TimelineKGQA example. Although TKG-Thinker is not trained on this dataset, it successfully adapts its tool usage strategy, verifies temporal consistency, and produces the correct answer. This result demonstrates that TKG-Thinker generalizes well to previously unseen temporal settings.